OpenStudy (anonymous):

Write the quadratic function in vertex form. y=x^2+4x+7

OpenStudy (anonymous):

now, do you know what the vertex form is?

OpenStudy (anonymous):

I did at one point in my life. Not so much anymore.

OpenStudy (anonymous):

Wait, no. It's y = a(x – h)^2 + k, right?

OpenStudy (igreen):

Yes, that's correct.

OpenStudy (igreen):

y = a(x – h)^2 + k y = 1(x - h)^2 + k h and k represent the vertex.

OpenStudy (anonymous):

Alright, I basically just need a step-by-step to let me know what I need to do and how to do it so I can figure out the problem.

OpenStudy (igreen):

h is the x-value k is the y-value

OpenStudy (igreen):

I just graph the equation and get the vertex, it takes to long to do it by hand.

OpenStudy (igreen):

The vertex is -2, 3

OpenStudy (igreen):

So we have: y = a(x - -2)^2 + 3 which equals: y = a(x + 2)^2 + 3

OpenStudy (igreen):

a = 1 (Hint: It's in the form of ax^2 + bx + c)

OpenStudy (igreen):

So our final answer is: y = 1(x + 2)^2 + 3

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

Alright! I got it now. I used to be able to do it pretty well but lost my skills a little. Thank you!

OpenStudy (igreen):

Np.